OpenStudy (anonymous):

A manufacturer of graphing calculators has determined that 15,000 calculators per week will be sold at a price of $96. At a price of $94, it is estimated that 15,740 calculators would be sold. (a) Determine a linear function that will predict the number of calculators y that would be sold at a given price x.

hero (hero):

Hint: apply slope formula 2nd Hint: apply point-slope formula

hero (hero):

3rd Hint: \[(x_1, y_1) = (15,000, 96)\]\[(x_2, y_2) = (15,740, 94)\]

OpenStudy (anonymous):

so 15000-15740/94-96

OpenStudy (anonymous):

well the answer I had in the end was 740/-3

hero (hero):

m = 740/-2 = -370

OpenStudy (anonymous):

ok thats where I think I was going wrong. Been a late night

OpenStudy (anonymous):

so then the next step would be -370x+b?

hero (hero):

Nevermind, we both made the same mistake

hero (hero):

It's correct except, we both had the wrong setup

OpenStudy (anonymous):

ok so its not -370?

hero (hero):

\[\frac{96 - 94}{15000 - 15740}\]

hero (hero):

No, it's not -370.... that's what I meant by "wrong setup"

OpenStudy (anonymous):

ok so 2/-740

hero (hero):

or - 1/370 always simplify to lowest terms

OpenStudy (anonymous):

ok and I take that to x and add it to what? the y1?

hero (hero):

No, that's our m = slope, value

hero (hero):

We need to apply this formula: \[y - y_1 = m(x - x_1)\]

OpenStudy (anonymous):

ok now I see where I am at

hero (hero):

\[y - 15000 = -\frac{(x-96)}{370}\]

hero (hero):

\[y - 15000 = \frac{96 - x}{370}\]

OpenStudy (anonymous):

then add the 1500 to both sides correct

hero (hero):

\[y = \frac{96 - x}{370} + 15000\]

hero (hero):

We can simplify that a little further

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

48-x/185?

hero (hero):

What did you do with the 15,000?

hero (hero):

That's the formula expressed in terms of one fraction

OpenStudy (anonymous):

I have not done anything yet but I will add it to the total minus x

OpenStudy (anonymous):

ok

hero (hero):

Hold on let me double check my work

hero (hero):

You should re-post the question. I must not have the right approach here because something is off

hero (hero):

Yeah, I think the x and y should be flipped

hero (hero):

It's the price that determines how many calculators will be sold. I apologize

hero (hero):

so m = -370

OpenStudy (anonymous):

ok

hero (hero):

Yeah, plus I mixed up the x and y while computing using the point-slope formula again. It's easy to get mixed up. I kept thinking the y value was 15,000 even though I was supposed to have 96 for the y value.. ugh

OpenStudy (anonymous):

ok got it its: -370x+50520

hero (hero):

Yes, that's it

OpenStudy (anonymous):

I submitted it and its correct. Thank you for your help

hero (hero):

Yes, it is correct

OpenStudy (anonymous):

thank you for your time and help